# SubTrack++ : Gradient Subspace Tracking for Scalable LLM Training

### v2

@@ Abstract
Training large language models (LLMs) is highly resource-intensive due to their massive number of parameters and the overhead of optimizer states. While recent work has aimed to reduce memory consumption, such efforts often entail trade-offs among memory efficiency, training time, and model performance. Yet, true democratization of LLMs requires simultaneous progress across all three dimensions. To this end, we propose SubTrack++ that leverages Grassmannian gradient subspace tracking combined with projection-aware optimizers, enabling Adam’s internal statistics to adapt to changes in the optimization subspace. Additionally, employing recovery scaling, a technique that restores information lost through low-rank projections, further enhances model performance. Our method demonstrates SOTA convergence by exploiting Grassmannian geometry and achieves lowest evaluation loss, outperforming the current SOTA while reducing pre-training wall time by 43% and maintaining the memory footprint on a 1B-parameter Llama model.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated state-of-the-art performance across a wide range of tasks and are rapidly growing in popularity. However, training and fine-tuning these models require significant resources, including extensive hardware and time, which limits their practicality for many applications and increases their environmental impact and carbon footprint. (Zhao et al., 2024; Jaiswal et al., 2024; Muhamed et al., 2024; Miles et al., 2024; Modoranu et al., 2024; Hao et al., 2024; Li et al., 2024).
Several techniques have been proposed to mitigate memory bottlenecks, such as gradient checkpointing (Chen et al., 2016) and memory offloading (Rajbhandari et al., 2020). LoRA (Hu et al., 2021) and other low-rank adaptation methods (Dettmers et al., 2024; Hu et al., 2021; Yaras et al., 2024; Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Miles et al., 2024) have gained popularity by optimizing a reduced set of parameters. Such approaches often assume that the full parameter space is inherently low-rank—an assumption that does not always hold and can lead to suboptimal performance. In addition, methods like BAdam (Luo et al., 2024) and Block-LLM (Ramesh et al., 2024), utilize block coordinate descent to optimize parameter subsets, achieving memory savings at the cost of partial parameter tuning and reduced accuracy.
@@
However, memory requirements extend beyond trainable parameters, with a significant portion consumed by optimizers for storing element-wise states (Zhao et al., 2024). To address this, recent efforts have focused on reducing the optimizer parameters while still targeting full parameter training (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022; Zhang et al., 2024; Modoranu et al., 2024; Zhao et al., 2024; Muhamed et al., 2024; Rajabi and Rambhatla, 2024a). Leveraging the fact that gradients evolve in a low-dimensional space during gradient descent (Gur-Ari et al., 2018; Schneider et al., 2024; Yaras et al., 2023) , GaLore (Zhao et al., 2024) reduces memory usage by projecting gradients into a low-rank subspace and periodically updating the rank-$r$ approximation via singular value decomposition (SVD)—an approach also used in more recent method, Fira (Chen et al., 2025). While SVD offers optimal low-rank approximation (Robert et al., 2025), performing it repeatedly can pose several challenges. First, it is computationally intensive, and alternatives like random projections (as in APOLLO (Zhu et al., 2025) and GoLore (He et al., 2025)) or approximation methods for estimating dominant singular values (as in LDAdam (Robert et al., 2025), Online Subspace Descent (Liang et al., 2024) and Rajabi and Rambhatla (2024b)) match or outperform SVD in practice. Moreover, SVD is sensitive to noise (Vaswani et al., 2018; He et al., 2025) and tends to degrade in late training stages when gradients are small, often hindering convergence (He et al., 2025).
@@
Geometry-based methods have demonstrated strong performance in various machine learning applications (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Rajabi and Rambhatla, 2024c; Blocker et al., 2023). In this context, using Grassmannian manifolds for subspace tracking has led to structurally embedded information, lower computational complexity, and improved performance (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Blocker et al., 2023; Chakraborty et al., 2017; Zhang and Balzano, 2016). Grassmannian is the set of $r$-dimensional subspaces in a d-dimensional space, and the ultimate goal of low-rank optimization is to find the proper subspace in this space. This line of work has demonstrated robustness and efficiency in high-dimensional, noisy environments (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Balzano et al., 2018; Chakraborty et al., 2017), outperforming online or robust Principle Component Analysis (PCA) techniques (Li, 2004; Balzano et al., 2018; Chakraborty et al., 2017). Their natural robustness against perturbations and strong theoretical guarantees make them particularly well-suited to tracking the evolving gradient subspaces encountered in LLM training. (Balzano, 2012; Balzano et al., 2011).
@@
Figure 2: Visualization of Grassmannian-based subspace tracking: Between subspace updates, gradients are projected onto a fixed subspace. The tangent vector $\nabla F$ is computed via the derivative of a loss function that measures the subspace estimation error. The subspace is then updated by moving along the corresponding geodesic, determined by $\nabla F$, on the Grassmannian manifold to minimize estimation error.
@@
To this end, we employ subspace tracking on Grassmannian geodesics (Figure 2), to develop a geometry-based, time- and memory-efficient method for tracking the gradient subspace. This way, instead of reconstructing low-rank approximations via expensive SVD or random projections, we can efficiently leverage previously computed subspaces and the estimation error to better adjust the projection. We also incorporate subspace shifts into Adam’s first and second momentum update rules, ensuring proper alignment with coordinate changes using a Projection-Aware Optimizer (Robert et al., 2025). Additionally, we recover and scale (Recovery Scaling) the gradient information lost during low-rank projection to boost performance by utilizing the scaling information of the low-rank optimizer Chen et al. (2025); Zhu et al. (2025).
@@
To summarize, SubTrack++ is a projection-aware geometry-based approach that supports full-parameter training and incorporates recovery scaling, offering superior time efficiency compared to SVD or PowerSGD methods (e.g., GaLore (Zhao et al., 2024), FiRA (Chen et al., 2025), LDAdam (Robert et al., 2025)), while maintaining GaLore’s memory footprint. It outperforms online PCA methods like Online Subspace Descent (Liang et al., 2024) in subspace tracking and achieves SOTA convergence and evaluation loss across all strong baselines; Comparison provided in Figure 1.
@@ 2 SubTrack++
Challenges of Low-Rank Optimization. Projecting gradients into a low-rank subspace reduces memory overhead and enables scalable LLM training, but it introduces important trade-offs. First, gradient subspaces can be unstable and require adaptive tracking; while SVD-based methods can capture these shifts (Zhao et al., 2024; Chen et al., 2025), they are computationally expensive. To address this, recent work has explored cheaper approximations and random projections (Robert et al., 2025; Liang et al., 2024; Zhu et al., 2025; He et al., 2025). Furthermore, optimizers like Adam assume a fixed coordinate system, so any changes in the subspace must be reflected in their internal states to preserve consistent momentum updates. Finally, low-rank projections inherently discard some gradient components, that recovering and utilizing these discarded signals can lead to improved performance (Chen et al., 2025; Zhu et al., 2025).
Overview. We propose SubTrack++ , a memory- and time-efficient method for full, simultaneous parameter tuning that embeds geometric insights into low-rank optimization. It improves efficiency and performance via three core components: 1) Subspace Tracking—Grassmannian-based updates that refine projections using estimation error and subspace history; 2) Projection-Aware Optimizer—adjusting Adam’s internal states to account for shifting subspaces; and 3) Recovery Scaling—rescaling discarded gradient components to reclaim lost information.
Subspace Tracking. In SubTrack++ , subspace estimation is framed as selecting a point on the Grassmannian—the manifold of all $d$-dimensional subspaces in an $n$-dimensional space (Bendokat et al., 2024)—which is the formal definition and goal of low-rank projections. This geometric perspective offers three key benefits: 1) By moving along Grassmann geodesics, it refines the subspace using prior subspace and estimation error, avoiding full reinitialization as done in (Zhao et al., 2024; Chen et al., 2025; He et al., 2025). 2) Updates rely on lightweight algebraic operations, minimizing time and memory costs. 3) Controlled subspace shifts improve robustness against noise and abrupt changes.
The initial subspace is computed using SVD as shown in (1). $G_{0}\in\mathbb{R}^{m\times n}$ is the gradient matrix at step 00; $U$, $S$, and $V$ are its SVD components, and $r$ is the specified rank.
@@
To track subspace drift, we adjust the core subspace every $k$ steps, defined as the subspace update interval, by minimizing its Euclidean distance to the current gradient. The objective is formalized in the cost function (2).
@@
The tangent vector $\nabla F$ provides the direction for adjusting the subspace by accounting for the error lying in the orthogonal complement. However, to minimize changes to the subspace, SubTrack++ first computes a rank-$1$ approximation of $\nabla F$, determined by its largest singular value and the corresponding singular vector obtained from its SVD, and represented as $\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ in the final equality of (4). This approximation is then used to update the subspace.
As shown by Edelman et al. (1998); Bendokat et al. (2024), from which we have included the necessary definitions and theorem in section 3, we can move along a Grassmannian geodesic guided by the computed tangent vector, taking a step of size $\eta$, as presented in (5).
$\small S_{t+1}(\eta)=(S_{t}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ \sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$ (5)
@@
        $G_{lr}=\arg\min_{A}\|(S_{t-1}A-G_{t})\|^{2}$, and $R=G_{t}-S_{t-1}G_{lr}$ $\nabla F=-2RG_{lr}^{\top}\approx\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ $S_{t}=(S_{t-1}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ \sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t-1}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$
@@
     $W_{t}\leftarrow W_{t-1}-\alpha\cdot\widehat{G}_{t}$ $-\alpha\cdot\Lambda_{t}$
@@
Projection-Aware Optimizer. In Adam, the first and second momentum update rule is shown in (6) and (7), respectively (reminder: $\widetilde{G}_{t}$ is the projection of gradient into low-rank subspace).
@@
Inspired by methods such as LDAdam (Robert et al., 2025), we emphasize the importance of updating optimizer states in a projection-aware manner to account for shifting subspaces; otherwise, misaligned projections can distort the optimizer’s performance. To address this, at each subspace update step, we modify Adam’s original update rules in (6) and (7), replacing them with projection-aware counterparts represented in (8) and (9), which reflect subspace changes into optimizer statistics (Robert et al., 2025). Further details regarding this projections can be found in Appendix C.
@@
Here $\alpha$ is the learning rate and $\phi_{t}(G_{t})$ is the scaling factor computed based on the low-rank gradient representation, $\widetilde{G}_{t}$, and the optimizer’s processed output, $\widetilde{G}_{t}^{O}$ as:
@@
As suggested by Fira (Chen et al., 2025), this additional term must be controlled using a gradient-clipping-like method to ensure training stability. Building on their observation, we also limit the growth rate of $\Lambda_{t}=\phi_{t}(G_{t})(G_{t}-S_{t}\widetilde{G}_{t})$ by a factor $\zeta$, and apply a correction whenever it exceeds this threshold as:
@@ 3 Theoretical Analysis
$\small S_{t+1}(\eta)=(S_{t}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ \sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$
@@ 4 Experiments and Results
To ensure fair comparison, we evaluated SubTrack++ across diverse models and datasets through pre-training and fine-tuning, measuring key metrics critical to LLM democratization.
Pre-Training Experiments. We pre-trained several Llama-based models on the C4 dataset, with results in Table 1. To ensure a fair comparison, we benchmarked against a diverse set of baselines. BAdam (Luo et al., 2024) is notably efficient in both memory and time, but its reliance on partial parameter updates leads to reduced performance. GaLore (Zhao et al., 2024) supports full-parameter training with low memory usage, though this comes at the cost of more than 150% increase in wall-time relative to AdamW in large models.
Table 1: We compare evaluation loss for pre-training Llama-based architectures on the C4 dataset over 10K iterations. SubTrack++ outperforms all other baselines in nearly every configuration. The best results are marked in bold, with the second-best performance underlined. LDAdam could not be run on the 7B configuration due to an out-of-memory error with our available resources.
@@
SubTrack++ (Ours) 3.43 3.24 3.29 4.51 4.49 4.62
As another relative baseline, Online Subspace Descent (Liang et al., 2024) uses Online PCA to track gradient subspace. LDAdam (Robert et al., 2025) and Fira (Chen et al., 2025) deliver strong performance at the cost of increasing wall-time even more than GaLore. Across these methods, SubTrack++ achieves SOTA results while using 14% less memory and running 43% faster than LDAdam, the most accurate baseline on 1B model. Hyperparameter details are provided in Appendix G, with a detailed wall-time and memory analysis in Appendix F.
@@
Figure 3: Ablation study comparing pure Grassmannian subspace tracking with incremental additions of the projection-aware optimizer and recovery scaling, leading to SubTrack++ . While Grassmannian tracking alone matches GaLore’s step-wise convergence, it significantly reduces wall-time, as shown in Figure LABEL:fig:modular.
Figure LABEL:fig:modular compares pre-training performance of a 1B-parameter Llama model across several baselines. SubTrack++ converges fastest in both steps and wall-time, emphasizing the power of geometry-aware optimization in boosting performance, reducing resource demands, and democratizing access to SOTA training methods.
@@
Ablation Study. We conducted an ablation study to evaluate the individual and combined effects of projection-aware optimizers and recovery scaling, integrated with Grassmannian subspace tracking on the 1B Llama model. We use the experimental settings outlined in Table 10. As shown in Figure 3, projection-aware optimizer and recovery scaling independently yield substantial improvements over pure Grassmannian subspace tracking, reducing loss from 6.53 to 5.43 and 5.28, respectively; while their combination, SubTrack++ , achieves 4.51 surpassing all baselines. Notably, these gains are achieved with minimal increases in runtime and memory.
@@
Fine-Tuning Experiments. RoBERTa-Base and RoBERTa-Large are fine-tuned on GLUE and SuperGLUE tasks; with the results presented in Table 4 and 5, respectively. More details and hyperparameters are provided in Appendix E.
Time and Space Complexity. Table 2 provides memory requirements of the optimizer states and the time complexity of the subspace update step considering an $m\times n$ gradient matrix with $m\leq n$. GaLore (Zhao et al., 2024) and Fira (Chen et al., 2025) periodically perform SVD to estimate the underlying subspace, while LDAdam (Robert et al., 2025) relies on the faster PowerSGD to update the subspace at every iteration. In contrast, SubTrack++ employs Grassmannian-based subspace tracking at the same frequency as GaLore and Fira. Comparing the time complexities of these methods, highlights why SubTrack++ is significantly more efficient than SVD-based methods. A breakdown of the subspace update time complexity for SubTrack++ is shown in Appendix D. Additionally, the memory required for storing optimizer states in SubTrack++ , is equivalent to GaLore and other baselines.

### v3

@@ Abstract
Training large language models (LLMs) is highly resource-intensive due to their massive number of parameters and the overhead of optimizer states. While recent work has aimed to reduce memory consumption, such efforts often entail trade-offs among memory efficiency, training time, and model performance. Yet, true democratization of LLMs requires simultaneous progress across all three dimensions. To this end, we propose SubTrack++ that leverages Grassmannian gradient subspace tracking combined with projection-aware optimizers, enabling Adam’s internal statistics to adapt to subspace changes. Additionally, employing recovery scaling, a technique that restores information lost through low-rank projections, further enhances model performance. Our method demonstrates SOTA convergence by exploiting Grassmannian geometry, reducing pre-training wall-time by up to 65% and fine-tuning time by 36% compared to existing SOTA methods, while maintaining the same memory footprint. Code is at https://github.com/criticalml-uw/SubTrack.
@@ 1 Introduction
LLMs have demonstrated state-of-the-art performance across a wide range of tasks and are rapidly growing in popularity. However, training and fine-tuning these models require significant resources, including extensive hardware and time, which limits their practicality for many applications and increases their environmental impact and carbon footprint. (Zhao et al., 2024; Jaiswal et al., 2024; Muhamed et al., 2024; Miles et al., 2024; Modoranu et al., 2024; Hao et al., 2024; Li et al., 2024).
Several techniques have been proposed to mitigate memory bottlenecks(Chen et al., 2016; Rajbhandari et al., 2020). LoRA (Hu et al., 2021) and other low-rank adaptation methods (Dettmers et al., 2024; Hu et al., 2021; Yaras et al., 2024; Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Miles et al., 2024) have gained popularity by optimizing a reduced set of parameters. Such approaches often assume a low-rank parameter space, which can lead to suboptimal performance. In addition, methods like BAdam (Luo et al., 2024) and Block-LLM (Ramesh et al., 2024), utilize block coordinate descent to optimize parameter subsets, achieving memory savings at the cost of reduced accuracy.
@@
However, memory requirements extend beyond trainable parameters, with a significant portion consumed by the optimizer’s states (Zhao et al., 2024). Recent efforts have focused on reducing this space while targeting full parameter training (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022; Zhang et al., 2024; Modoranu et al., 2024; Zhao et al., 2024; Muhamed et al., 2024). Leveraging the low-dimensional nature of gradients during gradient descent (Gur-Ari et al., 2018; Schneider et al., 2024; Yaras et al., 2023) , GaLore (Zhao et al., 2024) reduces memory usage by projecting gradients into a low-rank subspace and periodically updating this approximation via singular value decomposition (SVD). While SVD offers optimal low-rank approximation (Robert et al., 2025), it can pose several challenges. First, it is computationally intensive, and alternatives which use random projections (Zhu et al., 2025) or approximation methods to estimate dominant singular values (Robert et al., 2025; Liang et al., 2024), match or outperform SVD in practice. Moreover, SVD is sensitive to noise (Vaswani et al., 2018; He et al., 2025) and tends to degrade in late training stages when gradients are small, often hindering convergence (He et al., 2025).
@@
Geometry-based methods have shown strong performance in various machine learning applications (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Blocker et al., 2023); Grassmannian is the manifold of all subspaces of dimensions $r$ in a space of dimensions $d$, and using Grassmannian for subspace tracking has led to structurally embedded information, lower computational complexity, and improved performance (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Blocker et al., 2023; Chakraborty et al., 2017; Zhang and Balzano, 2016). This line of work has demonstrated robustness and efficiency in high-dimensional, noisy environments (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Balzano et al., 2018; Chakraborty et al., 2017). Their natural robustness against perturbations and strong theoretical guarantees make them particularly well-suited to tracking the evolving gradient subspaces encountered in LLM training.
@@
Figure 2: Visualization of Grassmannian subspace tracking: Between subspace updates, gradients are projected onto a fixed subspace. The tangent vector $\nabla F$ is computed via the derivative of a loss function, measuring the subspace estimation error. The subspace is then updated by moving along the corresponding geodesic, determined by $\nabla F$ to minimize estimation error.
@@
To this end, we employ subspace tracking on Grassmannian geodesics (Figure 2), to develop a geometry-based, time- and memory-efficient training method. This way, instead of reconstructing low-rank approximations via expensive SVD, we can efficiently leverage previously computed subspaces and the estimation error to better adjust the projection. We also incorporate subspace shifts into Adam’s first and second momentum update rules, ensuring proper alignment with coordinate changes using a Projection-Aware Optimizer (Robert et al., 2025). Additionally, we recover and scale (Recovery Scaling) the gradient information lost during low-rank projection by utilizing the scaling information of the low-rank, state-full optimizer Chen et al. (2025); Zhu et al. (2025).
@@
To summarize, SubTrack++ is a projection-aware geometry-based approach that supports full-parameter training and incorporates recovery scaling, offering superior time efficiency compared to SVD or PowerSGD methods (e.g., GaLore (Zhao et al., 2024; Chen et al., 2025; Robert et al., 2025), while maintaining GaLore’s memory footprint. It also outperforms online PCA subspace tracking methods (Liang et al., 2024) and achieves SOTA convergence and evaluation loss across all strong baselines; with comparison provided in Figure 1 and Figure 7.
@@ 2 SubTrack++
Challenges of Low-Rank Optimization. Projecting gradients into a low-rank subspace reduces memory footprint and enables scalable LLM training, but it introduces important trade-offs. First, gradient subspaces require adaptive tracking; while SVD-based methods can capture these shifts (Zhao et al., 2024; Chen et al., 2025), they are computationally expensive. To address this, recent work has explored cheaper approximations and random projections (Robert et al., 2025; Liang et al., 2024; Zhu et al., 2025; He et al., 2025). Furthermore, optimizers like Adam assume a fixed coordinate system, and subspace changes must be reflected in their internal states for a consistent momentum updates. Finally, low-rank projections inherently discard some gradient components, that recovering and utilizing these discarded signals can boost performance (Chen et al., 2025; Zhu et al., 2025).
Overview. SubTrack++ , a memory- and time-efficient method, embeds geometric insights into low-rank optimization, and improves efficiency and performance via three core components: 1) Grassmannian subspace tracking that refines projections via estimation error and subspace history; 2) projection-aware optimizer which adapts Adam’s state to account for evolving subspaces; and 3) recovery scaling that restores lost information by scaling discarded gradient components.
Subspace Tracking. In SubTrack++ , subspace estimation is framed as selecting a point on the Grassmannian, the manifold of all $d$-dimensional subspaces in an $n$-dimensional space (Bendokat et al., 2024). This geometric perspective offers three key benefits: 1) It refines the subspace using prior subspace and estimation error, avoiding full reinitialization as done in Zhao et al. (2024); Chen et al. (2025); He et al. (2025). 2) Updates rely on lightweight algebraic operations, minimizing time and memory costs. 3) Controlled subspace shifts improve robustness against noise and abrupt changes. The initial subspace is computed using SVD as shown in (1). $G_{0}\in\mathbb{R}^{m\times n}$ is the gradient matrix at step 0; $U$, $S$, and $V$ are its SVD components, and $r$ is the specified rank.
@@
To account for subspace drift, the core subspace is updated every $k$ steps (the subspace update interval) by minimizing the cost function (2), which measures its Euclidean distance to the current gradient.
@@
For optimizing the loss function (2), the subspace should be moved in the direction of $\nabla F$ to reduce the estimation error. However, to control subspace changes, SubTrack++ computes a rank-$1$ approximation of $\nabla F$, determined by its largest singular value and the corresponding singular vector obtained from its SVD, represented as $\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$. This approximation is then used for subspace update. As shown by Edelman et al. (1998); Bendokat et al. (2024), we can move along a Grassmannian geodesic guided by rank-1 estimation of $\nabla F$, with a step-size $\eta$, as presented in (5).
$\small S_{t+1}(\eta)=(S_{t}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ -\sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$ (5)
@@
   $G_{lr}=\arg\min_{A}\|(S_{t-1}A-G_{t})\|^{2}$, and $R=G_{t}-S_{t-1}G_{lr}$ $\nabla F=-2RG_{lr}^{\top}\approx\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$$S_{t}=(S_{t-1}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ -\sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t-1}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$
@@
  $W_{t}\leftarrow W_{t-1}-\alpha\cdot\widehat{G}_{t}$ $\alpha\cdot\Lambda_{t}$
@@
Projection-Aware Optimizer. In Adam, the first and second momentum update rules are as shown in (6) and (7), respectively (reminder: $\widetilde{G}_{t}$ is the projection of gradient into low-rank subspace).
@@
Inspired by methods such as LDAdam (Robert et al., 2025), we emphasize the importance of updating optimizer states in a projection-aware manner to account for shifting subspaces; otherwise, misaligned projections can distort the optimizer’s performance. To address this, at each subspace update step, we modify Adam’s original update rules in (6) and (7), replacing them with projection-aware counterparts represented in (8) and (9), which reflect subspace changes into optimizer statistics (Robert et al., 2025). Further details regarding these projections can be found in Appendix C.
@@
Here $\alpha$ is the learning rate and $\phi_{t}(G_{t})$ is the column-wise scaling factor computed based on the low-rank gradient representation, $\widetilde{G}_{t}$, and the optimizer’s processed output, $\widetilde{G}_{t}^{O}$ as:
@@
Following Fira’s observation (Chen et al., 2025), we employ a gradient-clipping-inspired mechanism to stabilize training. Specifically, we limit the growth rate of $\Lambda_{t}=\phi_{t}(G_{t})(G_{t}-S_{t}\widetilde{G}_{t})$ by a factor $\zeta$, and apply a correction whenever it exceeds this threshold as:
@@ 3 Theoretical Analysis
$\small S_{t+1}(\eta)=(S_{t}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ -\sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$
@@ 4 Experiments and Results
We evaluated SubTrack++ across diverse models and datasets through pre-training and fine-tuning, measuring key metrics critical to LLM democratization.
Table 1: We compare evaluation loss ($\downarrow$) for pre-training Llama-based architectures on the C4 dataset over 10k iterations. SubTrack++ outperforms all other baselines in nearly every configuration. The best results are marked in bold, with the second-best performance underlined. LDAdam could not be run on the 7B configuration due to an out-of-memory error with our available resources.
@@
SubTrack++ (Ours) 3.43 3.24 3.29 4.52 4.50 4.63
Pre-Training Experiments. We pre-trained several Llama-based models on the C4 dataset, with results in Table 1. To ensure a fair comparison, we benchmarked against a diverse set of baselines. While all compared methods aim for memory-efficient training, their architectural principles yield distinct computational and convergence trade-offs. BAdam (Luo et al., 2024) achieves strong memory and time efficiency, yet its partial parameter update strategy compromises final performance. GaLore (Zhao et al., 2024) constrains gradients to a low-rank subspace estimated via periodic SVD, an approach sensitive to noise and costly to compute, while discarding information residing in the orthogonal complement. Fira (Chen et al., 2025) introduces norm-based recovery scaling to mitigate this information loss, but its reliance on frequent SVD still leads to substantial wall-time overhead. LDAdam (Robert et al., 2025) replaces SVD with PowerSGD-based iterative updates and a projection-aware optimizer that synchronizes internal states with evolving subspaces, improving convergence and stability but incurring high per-step costs due to continual subspace updates. It also adds an extended error-feedback mechanism to compensate for both gradient and optimizer compression to target GaLore’s shortcomings. Online Subspace Descent (OSD) (Liang et al., 2024), our tracking-based baseline, further reduces complexity by employing Online PCA for subspace tracking.
SubTrack++ departs from these formulations by treating subspace evolution as a geometric tracking problem on the Grassmannian. It performs efficient rank-1 geodesic updates that reuse historical subspace information, inherently preserving consistency and avoiding the instability of discrete subspace resets. Simultaneously, it aligns Adam’s internal states through projection-aware optimizer and leverages recovery scaling to reintegrate the lost information. This unified approach yields a new training paradigm that retains the memory efficiency of low-rank methods, while achieving on-par runtime efficiency compared to fastest baselines like APOLLO (Zhu et al., 2025). SubTrack++ sets new state-of-the-art results across model scales, running 43% faster than LDAdam, the strongest prior baseline, on the 1B model, and 67% faster on the 3B model (see Table 6).
Figure LABEL:fig:modular demonstrates the pre-training of a 1B-parameter Llama model across several baselines. SubTrack++ has the fastest convergence in both training steps and wall-time, highlighting the effectiveness of geometry-aware optimization in improving performance and reducing resource consumption. To further assess its generalization in longer training regimes and larger models, we extend training to 100k steps and compare SubTrack++ with GaLore. As shown in Figure LABEL:fig:7b-long, SubTrack++ converges substantially faster, achieving an evaluation loss of 3.37 (a significant improvement over the 10k-step setting), while GaLore reaches 4.64 under the same conditions.
As shown in Table 1, SubTrack++ occasionally outperforms full-rank training. This effect may be attributed to the implicit regularization introduced by low-rank projections, which can enhance generalization in overparameterized models. Similar trends have been reported in other studies (Robert et al., 2025; Zhu et al., 2025; Chen et al., 2025). In addition, to further validate that convergence on the projected gradient reflects convergence of the original full-gradient, we measured both norms during Llama-1B pre-training. The full-gradient norm drops from 0.46 → 0.08, while the projected-gradient norm drops from 0.45 → 0.05, following nearly identical trajectories. This confirms that the optimization progress observed on the projected gradient accurately mirrors convergence in the original gradient space, consistent with prior low-rank gradient findings.
Hyperparameters of pre-training experiments are provided in Appendix E, with detailed runtime and memory reports in Appendix F.
(a) Loss vs. Training Steps.
(b) Loss vs. Relative Wall-Time.
@@
Figure 3: Ablation study comparing pure Grassmannian subspace tracking with incremental additions of the projection-aware optimizer and recovery scaling, leading to SubTrack++ . While Grassmannian tracking alone almost matches GaLore’s step-wise convergence (a), it significantly reduces wall-time (b).
@@
Ablation Studies. We conducted an ablation study to assess the individual and combined contributions of the projection-aware optimizer and recovery scaling, integrated with Grassmannian subspace tracking, on the 1B Llama model. Experimental settings are summarized in Table 4. As illustrated in Figure 3-b, Grassmannian subspace tracking alone substantially reduces wall-time compared to frequent SVD updates. Both the projection-aware optimizer and recovery scaling independently provide notable performance gains over baseline subspace tracking, lowering the loss from 6.53 to 5.43 and 5.28, respectively. Their combination, SubTrack++ , further improves the loss to 4.51, surpassing all baselines. Importantly, these improvements are achieved with only minimal increases in runtime and memory overhead, thanks to the efficiency of Grassmannian subspace tracking.
In addition, we conducted ablations on the subspace update rank and update frequency. As shown in Figure 4-a, more frequent updates can further improve performance, and the computational efficiency of SubTrack++ enables higher update frequencies with minimal overhead. However, excessively frequent updates may impede convergence. Figure 4-b shows that rank-1 updates achieve the best performance among all tested values. This suggests that making controlled, small adjustments to the underlying subspace helps maintain stability, while updating the subspace along the most informative direction prevents stagnation in a low-rank region and enhances generalization.
@@
Fine-Tuning Experiments. RoBERTa-Base and RoBERTa-Large are fine-tuned on GLUE (Wang et al., 2019) and SuperGLUE (Sarlin et al., 2020) tasks; with the results presented in Table 7 and 8, respectively. We also conducted supervised fine-tuning of the Llama-2-7B-chat-hf model for one epoch on the Alpaca (Taori et al., 2023) dataset. In this experiment, SubTrack++ achieved 36% lower wall-time and better performance compared to GaLore. The results are presented in Table 9. More details and hyperparameters are provided in Appendix G.
Time and Space Complexity. Table 2 provides memory requirements of the optimizer states and the time complexity of the subspace update step considering an $m\times n$ gradient matrix with $m\leq n$.